OpenStudy (anonymous):

Hey guys I'm having problems with derivatives if someone could explain the logx/lnx derivatives to me that would be great heres a few examples y = -2x^2(lnx)^4 or y = ln((x +1)/x)

OpenStudy (anonymous):

to start, do you know the derivative of \[\ln(x)\] ?

OpenStudy (anonymous):

\[d/dx [ \ln (x) ] = 1/x\]

OpenStudy (anonymous):

yes i just dont understand how it applies when you have things attached to x

OpenStudy (anonymous):

such as the (lnx)^4 or the ln(x +1/x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let's try \[\ln((x+1)/x)\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

have you used substitution rule before?

OpenStudy (anonymous):

That example is very simple, though the process might be tedious. Simply put a 1 over the expression then multiply by derivative of the expression

OpenStudy (anonymous):

No, i got that wrong, i was describing log.

OpenStudy (anonymous):

I'm contradicting myself, I think i told you right

OpenStudy (anonymous):

so, ln((x+1)/x)

OpenStudy (anonymous):

using the substitution rule: let's have \[u = (x+1)/x , du = -1/x^2\]

OpenStudy (anonymous):

du = =1/x^2 * dx

OpenStudy (anonymous):

yes, addmaster?

OpenStudy (anonymous):

ok sorry the substitution thing is confusing me

OpenStudy (anonymous):

ok, well, the concept is thus: essentially you can simply things by taking a difficult expression to immeditaely differentitate, like here ln((x+1)/x), and simplify the first steps a bit

OpenStudy (anonymous):

outright, [ln((x+1)/x)]' looks difficult, but [ln(u)]'

OpenStudy (anonymous):

that is like a calc identity

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

the trick is keeping track of this substitution, you'll need to find the derivtative of u. as above, u=(x+1)/x, while du = (-1/x^2)*dx

OpenStudy (anonymous):

ahhh ok so if you set it as a variable it complicates things less

OpenStudy (anonymous):

so we have our first equation: ln((x+1)/x) dx

OpenStudy (anonymous):

yeah, you do less internal variable manipulation

OpenStudy (anonymous):

alright got it sorry

OpenStudy (anonymous):

cool, you got it?

OpenStudy (anonymous):

yep that helped a lottt

OpenStudy (anonymous):

so, you want to solve it? i can stay to help if you need it...

OpenStudy (anonymous):

but just to be sure once I have the x + 1 over x I have to then use quotient rule...?

OpenStudy (anonymous):

so, u = (x+1)/x ... but that equals x/x + 1/x

OpenStudy (anonymous):

so u = 1 + 1/x

OpenStudy (anonymous):

du = (1/x) dx

OpenStudy (anonymous):

er, (1/x)'

OpenStudy (anonymous):

which is (\[x^-1\])'

OpenStudy (anonymous):

so, whats the derivative with respect to x of 1 + x^-1 ?

OpenStudy (anonymous):

-1x^-2...?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so, ln(u) du ---> but du = (-1/x^2)*dx

OpenStudy (anonymous):

ln(u)' --> 1/u, where u is (1 + 1/x) ....so you have \[1/[(1 + 1/x)]\]....then, you can't forget that du was -1/x^2 * dx, so when you now substitute back in for u so you have the original variable x, you have to multiply by (-1/x^2)

OpenStudy (anonymous):

\[\frac{1}{(1+1/x)} * \frac{-1}{(x^2)}\]

OpenStudy (anonymous):

which you can simplify

OpenStudy (anonymous):

to : -1/(x^2 + x)

OpenStudy (anonymous):

ok sorry my computers really crappy so i cant reply fast but i completely get it now thanks!

OpenStudy (anonymous):

no problem, hopefully now you can use the book examples much better too!

OpenStudy (anonymous):

also, here on OpenStudy, if someone gives you good help, you can award them with a medal here in the conversation

OpenStudy (anonymous):

so, ask more questions and maybe you can help some other math learners if you see a question you know to answer

OpenStudy (anonymous):

i'm off, but good luck with the calc ;)

OpenStudy (anonymous):

thank you very much